# In deep reinforcement learning, a pruned network is a good network

### v1

@@ Abstract
Recent work has shown that deep reinforcement learning agents have difficulty in effectively using their network parameters. We leverage prior insights into the advantages of sparse training techniques and demonstrate that gradual magnitude pruning enables agents to maximize parameter effectiveness. This results in networks that yield dramatic performance improvements over traditional networks and exhibit a type of “scaling law”, using only a small fraction of the full network parameters.
Abstract
Recent work has shown that deep reinforcement learning agents have difficulty in effectively using their network parameters. We leverage prior insights into the advantages of sparse training techniques and demonstrate that gradual magnitude pruning enables agents to maximize parameter effectiveness. This results in networks that yield dramatic performance improvements over traditional networks and exhibit a type of “scaling law”, using only a small fraction of the full network parameters.
@@ 1 Introduction
One of the most surprising findings of this last work is that applying the gradual magnitude pruning technique proposed by Zhu and Gupta (2017) on DQN (Mnih et al., 2015) with a ResNet backbone (as introduced in Impala (Espeholt et al., 2018)), results in a 50% performance improvement over the dense counterpart, with only 10% of the original parameters (see the bottom right panel of Figure 1 of Graesser et al. (2022)). Curiously, when the same pruning technique is applied to the original CNN architecture there are no performance improvements, but no degradation either.
Figure 1: Scaling network widths for ResNet architecture, for DQN and Rainbow with an Impala-based ResNet. We report the interquantile mean after 40 million environment steps, aggregated over 15 games with 5 seeds each; error bars indicate 95% stratified bootstrap confidence intervals.
@@
In this paper we explore gradual magnitude pruning as a general technique for improving the performance of RL agents. We demonstrate that in addition to improving the performance of standard network architectures, the gains increase proportionally with the size of the base network architecture. This last point is significant, as deep RL networks are known to struggle with scaling architectures (Ota et al., 2021; Farebrother et al., 2023; Taiga et al., 2023; Schwarzer et al., 2023). Our main contributions are as follows. We:
  • present gradual magnitude pruning as a general technique for maximizing parameter efficiency in RL;
  • demonstrate that networks trained with this technique produce stronger agents than their dense counterparts, and exhibit a type of “scaling law” as we scale the size of the initial network;
  • show this technique is generally useful across a varied set of agents and training regimes;
  • present in-depth analyses to better understand the reasons behind their benefits.
@@ 2 Related Work
Deep neural networks have been the driving factor behind many of the successful applications of reinforcement learning to real-world tasks. However, it has been historically difficult to scale these networks, in a manner similar to what has led to the “scaling laws” in supervised learning, without performance degradation; this is due in large part to exacerbated training instabilities that are endemic to reinforcement learning (Van Hasselt et al., 2018; Sinha et al., 2020; Ota et al., 2021). Recent works that have been able to do so successfully have had to rely on a number of targeted techniques and careful hyper-parameter selection (Farebrother et al., 2023; Taiga et al., 2023; Schwarzer et al., 2023; Ota et al., 2021; Ceron et al., 2023).
@@
Figure 2: Gradual magnitude pruning schedules used in our experiments, to a target sparsity of 95%, as specified in Equation 1.
@@
Previous studies (Schmitt et al., 2018; Zhang et al., 2019) have employed knowledge distillation with static data to mitigate instability, resulting in small, but dense, agents. Livne and Cohen (2020) introduced policy pruning and shrinking, utilizing iterative policy pruning similar to iterative magnitude pruning (Han et al., 2015), to obtain a sparse DRL agent. The exploration of the lottery ticket hypothesis in DRL was initially undertaken by Yu et al. (2019), and later Vischer et al. (2021) demonstrated that a sparse winning ticket can also be identified through behavior cloning. Sokar et al. (2021) proposed the use of structural evolution of network topology in DRL, achieving 50% sparsity with no performance degradation. Arnob et al. (2021) introduced single-shot pruning for offline Reinforcement Learning. Graesser et al. (2022) discovered that pruning often yields improved results, and dynamic sparse training methods, where the sparse topology changes throughout training (Mocanu et al., 2018; Evci et al., 2020), can significantly outperform static sparse training, where the sparse topology remains fixed throughout training. Tan et al. (2023) enhance the efficacy of dynamic sparse training through the introduction of a novel delayed multi-step temporal difference target mechanism and a dynamic-capacity replay buffer. Grooten et al. (2023) proposed an automatic noise filtering method, which uses the principles of dynamic sparse training for adjusting the network topology to focus on task-relevant features.
@@
Song et al. (2019) and Zhang et al. (2018) highlighted and the tendency of RL networks to overfit, while Nikishin et al. (2022) and Sokar et al. (2023) demonstrated the prevalence of plasticity loss in RL networks, leading to a decline in final performance.
Several strategies have been proposed to mitigate this, such as data augmentation (Yarats et al., 2021; Cetin et al., 2022), dropout (Gal and Ghahramani, 2016), and layer and batch normalization (Ba et al., 2016; Ioffe and Szegedy, 2015). Hiraoka et al. (2021) demonstrated the success of employing dropout and layer normalization in Soft Actor-Critic (Haarnoja et al., 2018), while Liu et al. (2020) identified that applying $\ell_{2}$ weight regularization on actors can enhance both on- and off-policy algorithms.
@@ 3 Background
The goal in Reinforcement Learning is to optimize the cumulative discounted return over a long horizon, and is typically formulated as a Markov decision process (MDP) $(\mathcal{X},\mathcal{A},P,r,\gamma)$. An MDP is comprised of a state space $\mathcal{X}$, an action space $\mathcal{A}$, a transition dynamics model $P:\mathcal{X}\times\mathcal{A}\rightarrow\Delta(\mathcal{X})$ (where $\Delta(X)$ is a distribution over a set $X$), a reward function $\mathcal{R}:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$, and a discount factor $\gamma\in[0,1)$. A policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$ formalizes an agent’s behaviour. For a policy $\pi$, $Q^{\pi}(\mathbf{x},\mathbf{a})$ represents the expected discounted reward achieved by taking action $\mathbf{a}$ in state $\mathbf{x}$ and subsequently following the policy $\pi$: $Q^{\pi}(\mathbf{x},\mathbf{a}):=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}\left(\mathbf{x}_{t},\mathbf{a}_{t}\right)|\mathbf{x}_{0}=x,\mathbf{a}_{0}=a\right]$. The optimal Q-function, denoted as $Q^{\star}(\mathbf{x},\mathbf{a})$, satisfies the Bellman recurrence:
$Q^{\star}(\mathbf{x},\mathbf{a})=\\ \mathbb{E}_{\mathbf{x}^{\prime}\sim P\left(\mathbf{x}^{\prime}\mid\mathbf{x},\mathbf{a}\right)}\left[\mathcal{R}(\mathbf{x},\mathbf{a})+\gamma\max_{\mathbf{a}^{\prime}}Q^{\star}\left(\mathbf{x}^{\prime},\mathbf{a}^{\prime}\right)\right]$
@@
Figure 5: Scaling replay ratio for Rainbow with the ResNet architecture with a width of $3$. Default replaly ratio is $0.25$. See Section 4.1 for training details.
@@
height 90pt depth 0 pt width 1.2 pt Refer to captionRefer to caption
@@
In supervised learning settings there is a broad interest in sparse training techniques, whereby only a subset of the full network parameters are trained/used (Gale et al., 2019). This is motivated by computational and space efficiency, as well as speed of inference. Zhu and Gupta (2017) proposed a polynomial schedule for gradually sparsifying a dense network over the course of training by pruning model parameters with low weight magnitudes.
@@
4 Pruning boosts deep RL
performance
@@ 4.1 Implementation details
For the base DQN and Rainbow agents we use the Jax implementations of the Dopamine library (Castro et al., 2018) with their default values. It is worth noting that Dopamine provides a “compact” version of the original Rainbow agent, using only multi-step updates, prioritized replay, and distributional RL. For all experiments we use the Impala architecture introduced by Espeholt et al. (2018), which is a 15-layer ResNet, unles otherwise specified. Our reasoning for this is not only because of the results from Graesser et al. (2022), but also due to a number of recent works demonstrating that this architecture results in generally improved performance (Schmidt and Schmied, 2021; Kumar et al., 2022; Schwarzer et al., 2023).
We use the JaxPruner (Lee et al., 2024) library for gradual magnitude pruning, as it already provides integration with Dopamine. We follow the schedule of Graesser et al. (2022): begin pruning the network 20% into training and stop at 80%, keeping the final sparse network fixed for the rest of training. Figure 2 illustrates the pruning schedules used in our experiments (for 95% sparsity).
We evaluate our agents on the Arcade Learning Environment (ALE) (Bellemare et al., 2013) on the same 15 games used by Graesser et al. (2022), chosen for their diversity [^1] Discussed in A.4 in Graesser et al. (2022).. For computational considerations, most experiments were conducted over 40 million frames (as opposed to the standard 200 million); in our investigations we found the qualitative differences between algorithms at 40 million frames to be mostly consistent with those at 100 million (e.g. see Figure 10).
We follow the guidelines outlined by Agarwal et al. (2021) for evaluation: each experiment was run with 5 independent seeds, and we report the human-normalized interquantile mean (IQM), aggregated across the 15 games, configurations, and seeds, along with 95% stratified bootstrap confidence intervals. All experiments were run on NVIDIA Tesla P100 GPUs, and each took approximately 2 days to complete.
@@ 4.2 Online RL
Our findings thus far suggest that the use of gradual magnitude pruning increases the parameter efficiency of these agents. If so, then these sparse networks should also be able to benefit from more gradient updates. The replay ratio, which is the number of gradient updates per environment step, measures exactly this; it is well-known that it is difficult to increase this value without performance degradation (Fedus et al., 2020; Nikishin et al., 2022; Schwarzer et al., 2023; D’Oro et al., 2023). In Figure 5 we can indeed confirm that the pruned architectures maintain a performance lead over the unpruned baseline even at high replay ratio values.
Figure 7: Evaluating performance on the full Atari 2600 suite. DQN (left) and Rainbow (right), both using the ResNet architecture with a width of 3. See Section 4.1 for training details.
@@ 4.3 Low data regime
Kaiser et al. (2020) introduced the Atari 100k benchmark to evaluate RL agents in a sample-constrained setting, allowing agents only 100k environment interactions. For this regime, Kostrikov et al. (2020) introduced DrQ, a variant of DQN which makes use of data augmentation; the hyperparameters for this agent were further optimized by Agarwal et al. (2021) in DrQ($\epsilon$). Similarly, Van Hasselt et al. (2019) introduced Data-Efficient Rainbow (DER), which optimized the hyperparameters of Rainbow (Hessel et al., 2018) for this low data regime.
@@
Figure 8: Performance of DrQ($\epsilon$) (left) and DER (right) when trained for 40M frames. Both agents use a ResNet architecture with a width of 3. See Section 4.1 for training details.
@@ 4.4 Offline RL
Figure 10: Impact of varying pruning schedules, for DQN with an Impala-based ResNet (width = 3).
@@ 4.5 Actor-Critic methods
Our investigation thus far has focused on value-based methods. Here we investigate if gradual magnitude pruning can yield performance gains for Soft Actor Critic (SAC; Haarnoja et al., 2018), a popular policy-gradient algorithm. We evaluated SAC on five continuous control environments from the MuJoCo suite (Todorov et al., 2012), using 10 independent seeds for each. In Figure 10 we present the results for Walker2d-v2 and Ant-v2, where we see the advantages of gradual magnitude pruning persist; in the remaining three environments (see Appendix E) there are no real observable gains from pruning.

### v3

# a pruned network is a good network

@@ Abstract
Recent work has shown that deep reinforcement learning agents have difficulty in effectively using their network parameters. We leverage prior insights into the advantages of sparse training techniques and demonstrate that gradual magnitude pruning enables value-based agents to maximize parameter effectiveness. This results in networks that yield dramatic performance improvements over traditional networks, using only a small fraction of the full network parameters. Our code is publicly available, see Appendix A for details.
Machine Learning, ICML
@@ 1 Introduction
One of the most surprising findings of this last work is that applying the gradual magnitude pruning technique proposed by Zhu & Gupta (2017) on DQN (Mnih et al., 2015) with a ResNet backbone (as introduced in Impala (Espeholt et al., 2018)), results in a 50% performance improvement over the dense counterpart, with only 10% of the original parameters (see the bottom right panel of Figure 1 of Graesser et al. (2022)). Curiously, when the same pruning technique is applied to the original CNN architecture there are no performance improvements, but no degradation either.
Figure 1: Scaling network widths for ResNet architecture, for DQN and Rainbow with an Impala-based ResNet (Espeholt et al., 2018). We report the interquantile mean after 40 million environment steps, aggregated over 15 games with 5 seeds each; error bars indicate 95% stratified bootstrap confidence intervals. Replay ratio is fixed to the standard $0.25$. The default network is Dense, which we indicate with a blue color in all the plots, for clarity.
@@
In this paper we explore gradual magnitude pruning as a general technique for improving the performance of RL agents. We demonstrate that in addition to improving the performance of standard network architectures for value-based agents, the gains increase proportionally with the size of the base network architecture. This last point is significant, as deep RL networks are known to struggle with scaling architectures (Ota et al., 2021; Farebrother et al., 2023; Taiga et al., 2023; Schwarzer et al., 2023).
Our main contributions are as follows. We: present gradual magnitude pruning as a general technique for maximizing parameter efficiency in value-based RL; demonstrate that networks trained with this technique produce stronger agents than their dense counterparts, and continue improving as we scale the size of the initial network; investigate this technique across a varied set of agents and training regimes, including actor-critic methods; present in-depth analyses to better understand the reasons behind their benefits.
@@ 2 Related Work
Deep neural networks have been the driving factor behind many of the successful applications of reinforcement learning to real-world tasks. However, it has been historically difficult to scale these networks, in a manner similar to what has led to the “scaling laws” in supervised learning, without performance degradation; this is due in large part to exacerbated training instabilities that are endemic to reinforcement learning (Van Hasselt et al., 2018; Sinha et al., 2020; Ota et al., 2021). Recent works that have been able to do so successfully have had to rely on a number of targeted techniques and careful hyper-parameter selection (Farebrother et al., 2023; Taiga et al., 2023; Schwarzer et al., 2023; Ceron et al., 2023).
@@
Ceron et al. (2024) demonstrate that while parameter scaling with convolutional networks hurts single-task RL performance on Atari, incorporating Mixture-of-Expert (MoE) modules in such networks improves performance. Farebrother et al. (2024) demonstrate that value functions trained with categorical cross-entropy significantly improves performance and scalability in a variety of domains.
@@
Previous studies (Schmitt et al., 2018; Zhang et al., 2019) have employed knowledge distillation with static data to mitigate instability, resulting in small, but dense, agents. Livne & Cohen (2020) introduced policy pruning and shrinking, utilizing iterative policy pruning similar to iterative magnitude pruning (Han et al., 2015), to obtain a sparse DRL agent. The exploration of the lottery ticket hypothesis in DRL was initially undertaken by Yu et al. (2019), and later Vischer et al. (2021) demonstrated that a sparse winning ticket can also be identified through behavior cloning. Sokar et al. (2021) proposed the use of structural evolution of network topology in DRL, achieving 50% sparsity with no performance degradation. Arnob et al. (2021) introduced single-shot pruning for offline Reinforcement Learning.
Graesser et al. (2022) discovered that pruning often yields improved results, and dynamic sparse training methods, where the sparse topology changes throughout training (Mocanu et al., 2018; Evci et al., 2020), can significantly outperform static sparse training, where the sparse topology remains fixed throughout training. Tan et al. (2023) enhance the efficacy of dynamic sparse training through the introduction of a novel delayed multi-step temporal difference target mechanism and a dynamic-capacity replay buffer. Grooten et al. (2023) proposed an automatic noise filtering method, which uses the principles of dynamic sparse training for adjusting the network topology to focus on task-relevant features.
@@
Song et al. (2019) and Zhang et al. (2018) highlighted and the tendency of RL networks to overfit, while Nikishin et al. (2022) and Sokar et al. (2023) demonstrated the prevalence of plasticity loss in RL networks, leading to a decline in final performance. Several strategies have been proposed to mitigate this, such as data augmentation (Yarats et al., 2021; Cetin et al., 2022), dropout (Gal & Ghahramani, 2016), and layer and batch normalization (Ba et al., 2016; Ioffe & Szegedy, 2015). Hiraoka et al. (2021) demonstrated the success of employing dropout and layer normalization in Soft Actor-Critic (Haarnoja et al., 2018), while Liu et al. (2020) identified that applying $\ell_{2}$ weight regularization on actors can enhance both on- and off-policy algorithms.
@@ 3 Background
Figure 2: Gradual magnitude pruning schedules used in our experiments, to a target sparsity of 95%, as specified in Equation 1. Impact of varying pruning schedules, see Figure 10.
@@
The goal in Reinforcement Learning is to optimize the cumulative discounted return over a long horizon, and is typically formulated as a Markov decision process (MDP) $(\mathcal{X},\mathcal{A},P,r,\gamma)$. An MDP is comprised of a state space $\mathcal{X}$, an action space $\mathcal{A}$, a transition dynamics model $P:\mathcal{X}\times\mathcal{A}\rightarrow\Delta(\mathcal{X})$ (where $\Delta(X)$ is a distribution over a set $X$), a reward function $\mathcal{R}:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$, and a discount factor $\gamma\in[0,1)$. A policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$ formalizes an agent’s behaviour.
For a policy $\pi$, $Q^{\pi}(\mathbf{x},\mathbf{a})$ represents the expected discounted reward achieved by taking action $\mathbf{a}$ in state $\mathbf{x}$ and subsequently following the policy $\pi$: $Q^{\pi}(\mathbf{x},\mathbf{a}):=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}\left(\mathbf{x}_{t},\mathbf{a}_{t}\right)|\mathbf{x}_{0}=x,\mathbf{a}_{0}=a\right]$. The optimal Q-function, denoted as $Q^{\star}(\mathbf{x},\mathbf{a})$, satisfies the Bellman recurrence
$Q^{\star}(\mathbf{x},\mathbf{a})=\mathbb{E}_{\mathbf{x}^{\prime}\sim P\left(\mathbf{x}^{\prime}\mid\mathbf{x},\mathbf{a}\right)}\left[\mathcal{R}(\mathbf{x},\mathbf{a})+\gamma\max_{\mathbf{a}^{\prime}}Q^{\star}\left(\mathbf{x}^{\prime},\mathbf{a}^{\prime}\right)\right]$.
@@
In supervised learning settings there is a broad interest in sparse training techniques, whereby only a subset of the full network parameters are trained/used (Gale et al., 2019). This is motivated by computational and space efficiency, as well as speed of inference. Zhu & Gupta (2017) proposed a polynomial schedule for gradually sparsifying a dense network over the course of training by pruning model parameters with low weight magnitudes.
@@
4 Pruning can boost deep RL performance
@@ 4.1 Implementation details
For the base DQN and Rainbow agents we use the Jax implementations of the Dopamine library [^1] Dopamine code available at https://github.com/google/dopamine. (Castro et al., 2018) with their default values. It is worth noting that Dopamine provides a “compact” version of the original Rainbow agent, using only multi-step updates, prioritized replay, and distributional RL. For all experiments we use the Impala architecture introduced by Espeholt et al. (2018), which is a 15-layer ResNet, unless otherwise specified. Our reasoning for this is not only because of the results from Graesser et al. (2022), but also due to a number of recent works demonstrating that this architecture results in generally improved performance (Schmidt & Schmied, 2021; Kumar et al., 2022; Schwarzer et al., 2023).
We use the JaxPruner [^2] JaxPruner code available at https://github.com/google-research/jaxpruner. (Lee et al., 2024) library for gradual magnitude pruning, as it already provides integration with Dopamine. We follow the schedule of Graesser et al. (2022): begin pruning the network 20% into training and stop at 80%, keeping the final sparse network fixed for the rest of training. Figure 2 illustrates the pruning schedules used in our experiments (for 95% sparsity). We evaluate our agents on the Arcade Learning Environment (ALE) (Bellemare et al., 2013) on the same 15 games used by Graesser et al. (2022), chosen for their diversity [^3] Discussed in A.4 in Graesser et al. (2022).. For computational considerations, most experiments were conducted over 40 million frames (as opposed to the standard 200 million); in our investigations we found the qualitative differences between algorithms at 40 million frames to be mostly consistent with those at 100 million (e.g. see Figure 10).
We follow the guidelines outlined by Agarwal et al. (2021) for evaluation: each experiment was run with 5 independent seeds, and we report the human-normalized interquantile mean (IQM), aggregated across the 15 games, configurations, and seeds, along with 95% stratified bootstrap confidence intervals. All experiments were run on NVIDIA Tesla P100 GPUs, and each took approximately 2 days to complete. All hyper-parameters are listed in Appendix F.
Figure 5: Scaling replay ratio for Rainbow with the ResNet architecture with a width multiplier of $3$. Default replaly ratio is $0.25$. See Section 4.1 for training details.
@@ 4.2 Online RL
Figure 6: Evaluating performance on the full Atari 2600 suite. DQN (left) and Rainbow (right), both using the ResNet architecture with a width of $3$. We report IQM performance with error bars indicating 95% confidence interval. See Section 4.1 for training details.
@@
When evaluating on more modern value-based agents, specifically IQN (Dabney et al., 2018) and Munchausen-IQN (Vieillard et al., 2020), we observe the same advantages arising from pruning (see Section G.2).
Our findings thus far suggest that the use of gradual magnitude pruning increases the parameter efficiency of these agents. If so, then these sparse networks should also be able to benefit from more gradient updates. The replay ratio [^4] In the hyperparameters established in (Mnih et al., 2015), the policy is updated every 4 environment steps collected, resulting in a replay ratio of 0.25., which is the number of gradient updates per environment step, measures exactly this; it is well-known that it is difficult to increase this value without performance degradation (Fedus et al., 2020; Nikishin et al., 2022; Schwarzer et al., 2023; D’Oro et al., 2023).
In Figure 5 we can indeed confirm that the pruned architectures maintain a performance lead over the unpruned baseline even at high replay ratio values. The sharper rate of decline with pruning may suggest that the pruning schedule needs to be re-tuned for these settings.
@@ 4.3 Low data regime
Kaiser et al. (2020) introduced the Atari 100k benchmark to evaluate RL agents in a sample-constrained setting, allowing agents only 100k [^5] Here, 100k refers to agent steps, or 400k environment frames, due to skipping frames in the standard training setup. environment interactions. For this regime, Kostrikov et al. (2020) introduced DrQ, a variant of DQN which makes use of data augmentation; the hyperparameters for this agent were further optimized by Agarwal et al. (2021) in DrQ($\epsilon$). Similarly, Van Hasselt et al. (2019) introduced Data-Efficient Rainbow (DER), which optimized the hyperparameters of Rainbow (Hessel et al., 2018) for this low data regime.
@@
Figure 7: Performance of DrQ($\epsilon$) (left) and DER (right) when trained for 40M frames. Both agents use a ResNet architecture with a width multiplier of $3$. See Section 4.1 for training details.
@@ 4.4 Offline RL
height 98pt depth 0 pt width 1 pt Refer to captionRefer to caption
Figure 10: Impact of varying pruning schedules, for DQN with an Impala-based ResNet with a width multiplier of $3$.
@@ 4.5 Actor-Critic methods
Our investigation thus far has focused on value-based methods. Here we investigate if gradual magnitude pruning can yield performance gains for Soft Actor Critic (SAC; Haarnoja et al., 2018), a popular policy-gradient algorithm. We evaluated SAC on five continuous control environments from the MuJoCo suite (Todorov et al., 2012), using 10 independent seeds for each. In Figure 10 we present the results for Walker2d-v2 and Ant-v2, where we see the advantages of gradual magnitude pruning persist; in the remaining three environments (see Appendix E) there are no real observable gains from pruning. In Section G.1 we see a similar trend with PPO (Schulman et al., 2017).
@@ 4.6 Stability of the pruned network
4.7 Learning rate and Batch size scaling
The default learning rate or batch size may not be optimal for large neural networks. The default learning for DQN is $6.25\times 10^{-5}$. We run experiments with a learning rate divided by the width scale factor (so $2.08\times 10^{-5}$ for a width factor of $3$, and $1.25\times 10^{-5}$ for a width factor of $5$). While these learning rates do improve the performance of the baseline, it is still surpassed by pruning (see Figure 28). We observe a similar trend when evaluating different batch size values. The default batch size is 32 (for all value based agents used in this paper), and we ran experiments with batch sizes of 16 and 64. In all cases, pruning maintains its strong advantage (see Figure 29). These results are consistent with the thesis that pruning can serve as a drop-in mechanism for increasing agent performance.